# Learning in Structured Stackelberg Games

### v2

@@ 4.4 Upper bound
The idea here is to bound the probability that there exists a hypothesis yielding maximum utility for the leader over the sample, but is $\gamma$-far from the optimal hypothesis, by the probability that this hypothesis has poor performance on another sample. More concretely, it can be shown that $\mathscr{D}^{n}(Q_{\epsilon,\gamma})\leq 2\mathscr{D}^{2n}(R_{\epsilon,\gamma})$ for $n\geq c/\epsilon$, for some constant $c$. To see this, we first write
$\mathscr{D}^{2n}(R_{\epsilon,\gamma})=\int_{Q_{\epsilon,\gamma}}\mathscr{D}^{n}(\{S\in\pazocal{M}^{n}:\exists h\in\pazocal{H}:\hat{\text{er}}_{N}(h)=0,\hat{\text{er}}_{S,\gamma}(h)\geq\epsilon/2\})d\mathscr{D}^{n}(N)$
For some $N\in Q_{\epsilon,\gamma}$ consider some $h_{N}\in\pazocal{H}$ that satisfies $\hat{\text{er}}_{N,\gamma}(h_{N})=0$ and $\text{er}_{\mathscr{D},\gamma}(h_{N})>\epsilon$. We show that $\mathscr{D}(\hat{\text{er}}_{S,\gamma}(h_{S})\geq\epsilon/2)\geq 1/2$, from which the claim follows straightforwardly. We know that $\text{er}_{\mathscr{D},\gamma}(h_{N})>\epsilon$, and therefore $n\cdot\hat{\text{er}}_{S,\gamma}(h_{N})\leq n\cdot\hat{\text{er}}_{S}(h_{N})$ follows a binomial distribution with probability of success at least $\epsilon$. We note that the inequality above follows since mispredicting the follower type is a necessary condition for suboptimal utility. Using a Chernoff bound, we can show that
$\mathscr{D}^{n}(S:\hat{\text{er}}_{S,\gamma}(h_{N})<\epsilon/2)\leq e^{-n\epsilon/8}$
@@
The next step is to use a random swap argument to upper bound $\mathscr{D}^{2n}(R_{\epsilon,\gamma})$ with something that is independent of $\mathscr{D}$. We let $\Gamma_{n}$ denote the set of all permutations on $\{1,\dots,2n\}$ that potentially swap the $i$’th and $n+i$’th elements for all $i\in[n]$. So for all $\sigma\in\Gamma_{n}$ and $i\in[n]$, we have that either $\sigma(i)=i,\sigma(n+i)=n+i$ or $\sigma(i)=n+i,\sigma(n+i)=i$. From Lemma 3 in Attias et al. [3] (cf. also [2]), we get that for some $\pazocal{M}=\pazocal{Z}\times[K]$ and distribution $\mathscr{D}$ on $V$:
$\mathscr{D}^{2n}(R_{\epsilon,\gamma})=\mathop{\mathbb{E}}_{Z\sim\mathscr{D}^{2n}}\mathbb{P}_{\sigma\sim\text{Unif}(\Gamma_{n})}[\sigma(V)\in R_{\epsilon,\gamma}]\leq\max_{Z\in\pazocal{Z}^{2n}}\mathbb{P}_{\sigma\sim\text{Unif}(\Gamma_{n})}[\sigma(V)\in R_{\epsilon,\gamma}].$
The last step of the proof, requires reducing to learning a partial binary hypothesis class and using a disambiguation argument to turn the partial class into a total one, as in Attias et al. [3]. Concretely, we start by fixing a set $V\in\pazocal{M}^{2n}$, where $v_{i}=(\mathbf{z}_{i},h^{*}(\mathbf{z}_{i}))$ and a set $S=\{\mathbf{z}_{1},\dots,\mathbf{z}_{2n}\}$. For the projection, $\pazocal{H}(S)$ of the hypothesis class on the set, one can define the partial binary hypothesis class
@@
We know that if there exists such an $h$ that is correct on the first $n$ points but $\gamma$-far on at least $\epsilon n/2$ of the last $n$ points, there exists a function in $\tilde{\pazocal{H}}^{\prime}$, that is 0 on the first $n$ points and 111 on the points on which $h$ is $\gamma$-far. We also know that if for some $\sigma\in\Gamma_{n}$, $\tilde{h}^{\prime}\in\pazocal{H}^{\prime}$ is a witness that $\sigma(V)\in R_{\epsilon,\gamma}$, then at least one of $\tilde{h}^{\prime}(i),\tilde{h}^{\prime}(i)$ must be 0, and at least $n\epsilon/2$ elements must be non-zero. Thus for a permutation drawn uniformly at random, the probability that all 1’s are on the last n elements is at most $2^{-n\epsilon/2}$. Also since $|\tilde{\pazocal{H}}^{\prime}|$ is bounded we an use a union bound that yields $\mathbb{P}_{\sigma\sim\text{Unif}(\Gamma_{n})}[\sigma(V)\in R_{\epsilon,\gamma}]\leq n^{O(\textnormal{SGdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})\log(n))}2^{-n\epsilon/2}$. Given that $V\in\pazocal{M}^{2n}$, we can take a $\max$ over all $V\in\pazocal{M}^{2n}$, which gives the desired result. ∎
Remark 4.16.
It is worth noting that even when there is only one context (i.e. $\pazocal{Z}$ is a singleton), our results in this section are incomparable with those in Section˜4.1. This is because in this section, the goal of the leader is to learn given a deterministic mapping from contexts to follower types. So, if the contexts are the same between all examples in the training set, the follower types corresponding to these examples will also be the same. In contrast, the context-free setting we consider in Section˜4.1 allows for a distribution over follower types. Thus the two bounds are generally incomparable.
5 Computational Complexity Considerations
Computational complexity has been a longstanding concern in the analysis of Stackelberg games. In the simplest setting, where a leader interacts with a single known follower type Conitzer and Sandholm [20] have shown that computing the optimal mixed strategy for the leader can be achieved in polynomial time using linear programming. However, in more complex Stackelberg games such as Bayesian Stackelberg games, Stackelberg security games, and even 3-player Stackelberg games computing the leader’s optimal strategy has been shown to be NP-hard [20, 28]. Therefore computational considerations have so far been largely ignored. In our setting, there are several interesting cases of structured Stackelberg games for which the leader’s optimal strategy can be learned in polynomial time with provable (albeit not optimal) regret guarantees.
Observation.
In a structured Stackelberg game with hypothesis class $\pazocal{H}$ the leader’s optimal policy can be learned in polynomial time if $\pazocal{H}$ is efficiently online learnable. We give a high-level overview of this approach. First, we consider a learning algorithm for the hypothesis class $\pazocal{H}$. The leader can use this algorithm to predict the follower’s type and based on that compute the optimal mixed strategy against that type as if her prediction were correct. As shown by Conitzer and Sandholm [20], for a fixed and known follower type, computing the leader’s optimal mixed strategy can be achieved in polynomial time using linear programming. The learner will make a finite number of follower type mispredictions before finding $h^{*}$, and at this point will play optimally against every follower type that appears in every subsequent round. Therefore, for any hypothesis class in the realizable setting that is efficiently online- (and therefore also PAC-) learnable we can establish a computationally efficient algorithm for learning the leader’s optimal policy. We now present one such example of a hypothesis class that can be learned efficiently when used to predict the followers’ types.
Example.
@@ 5 Computational Complexity Considerations
Theorem 5.1.
@@
We present the full proof of Theorem˜5.1 and the corresponding algorithm in Section˜A.5.
6 Conclusion and Future Directions
We initiate the study of structured Stackelberg games, where contextual information is predictive of the follower’s type. Our main results include a complete characterization of instance-optimal regret bounds in the online setting and PAC sample complexity guarantees in the distributional setting. In this section, we discuss concrete open problems and future directions.
Computationally efficient algorithms.
An interesting avenue for future research is to consider the computational complexity of structured Stackelberg game algorithms. Although we have shown how to achieve optimal regret bounds, our proposed algorithm is not computationally efficient. As we discuss in Section˜5, our model opens the door for efficient, yet so far sub-optimal learning algorithms in certain settings. It would be interesting to further explore whether the structure introduced in our work can lead to both optimal and efficient algorithms or if there is an underlying tradeoff between computational complexity and tight regret guarantees.
Regret guarantees for the agnostic setting.
Our work focuses on structured Stackelberg games in the realizable setting, where the leader is given a class of hypotheses with the promise that it includes the optimal predictor. The agnostic learning setting is a natural relaxation of this assumption. While we provide some results (cf. Section˜B.1.4) for this setting, our bounds are not instance optimal. Providing a complete characterization for the agnostic case, analogously to the realizable one we study here, remains an avenue for future research. It is plausible that an entirely new complexity measure will be required here, which would show a clear separation between our setting and multiclass classification, where the Littlestone dimension characterizes both the realizable and the agnostic setting.
Beyond worst-case analysis frameworks.
Another direction for future research is to analyze structured Stackelberg games in the beyond worst-case analysis framework. It has recently been shown that smooth online learning is as easy as distributional learning, with regret bounds scaling with the VC dimension of the hypothesis class. Extending these findings to multiclass learning and structured Stackelberg games might lead to improved results both from a statistical and computational point of view.
Appendix A Details and Omitted Proofs
A.1 A technical subtlety
An important component of our analysis is that the leader must be able to compute the optimal strategy for a given context and follower type. For this, we state that for any context $\mathbf{z}\in\pazocal{Z}$ the leader incurs negligible regret by restricting herself to a subset $\mathscr{E}_{\mathbf{z}}\subseteq\Delta(\pazocal{A})$ which is finite and computable. To formally state our result we present the notion of a contextual best-response region, which was first introduced in [27] and which we repeat here for completeness.
Definition A.1 (Contextual Follower Best-Response Region).
Definition A.2 ($\delta$-approximate extreme points).
[…]
Notice that for sufficiently small $\delta$ i.e. $\delta=\mathscr{O}(\frac{1}{T})$, we get that the additional regret of the leader by choosing actions over $\mathscr{E}^{(i)}$ rather than $X(f^{(i)},a_{f})$ is negligible. Therefore, we use the shorthand $\mathscr{E}^{(i)}:=\mathscr{E}^{(i)}(\delta)$ for the rest of the section. The following lemma from [27] formalizes this notion.
Lemma A.3.
For any sequence of followers $f_{1},\ldots f_{T}$ and any leader policy $\pi$, there exists a policy $\pi^{(\mathscr{E})}:\pazocal{Z}\rightarrow\cup_{\mathbf{z}\in\pazocal{Z}}\mathscr{E}_{\mathbf{z}}$ that, when given context $\mathbf{z}$, plays a mixed strategy in $\mathscr{E}_{\mathbf{z}}$ and guarantees that $\sum_{t=1}^{T}u(\mathbf{z}_{t},\pi(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi(\mathbf{z}_{t})))-u(\mathbf{z}_{t},\pi^{(\mathscr{E})}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi^{(\mathscr{E})}(\mathbf{z}_{t})))\leq 1$. Moreover, the same result holds in expectation over any distribution over follower types.
Given Lemma˜A.3 and for ease of notation, in this paper we assume that the leader will always set the parameter $\delta$, such that it does not affect our regret bounds.
A.2 Proof of Theorem˜3.5
See 3.5

### v3

@@ Appendix C Details and proofs from Section˜4
The idea here is to bound the probability that there exists a hypothesis yielding maximum utility for the leader over the sample, but is $\gamma$-far from the optimal hypothesis, by the probability that this hypothesis has poor performance on another sample. More concretely, it can be shown that $\mathcal{D}^{n}(Q_{\epsilon,\gamma})\leq 2\mathcal{D}^{2n}(R_{\epsilon,\gamma})$ for $n\geq c/\epsilon$, for some constant $c$. To see this, we first write
$\mathcal{D}^{2n}(R_{\epsilon,\gamma})=\int_{Q_{\epsilon,\gamma}}\mathcal{D}^{n}(\{S\in\pazocal{M}^{n}:\exists h\in\pazocal{H}:\hat{\text{er}}_{N}(h)=0,\hat{\text{er}}_{S,\gamma}(h)\geq\epsilon/2\})d\mathcal{D}^{n}(N)$
For some $N\in Q_{\epsilon,\gamma}$ consider some $h_{N}\in\pazocal{H}$ that satisfies $\hat{\text{er}}_{N,\gamma}(h_{N})=0$ and $\text{er}_{\mathcal{D},\gamma}(h_{N})>\epsilon$. We show that $\mathcal{D}(\hat{\text{er}}_{S,\gamma}(h_{S})\geq\epsilon/2)\geq 1/2$, from which the claim follows straightforwardly. We know that $\text{er}_{\mathcal{D},\gamma}(h_{N})>\epsilon$, and therefore $n\cdot\hat{\text{er}}_{S,\gamma}(h_{N})\leq n\cdot\hat{\text{er}}_{S}(h_{N})$ follows a binomial distribution with probability of success at least $\epsilon$. We note that the inequality above follows since mispredicting the follower type is a necessary condition for suboptimal utility. Using a Chernoff bound, we can show that
$\mathcal{D}^{n}(S:\hat{\text{er}}_{S,\gamma}(h_{N})<\epsilon/2)\leq e^{-n\epsilon/8}$
@@
The next step is to use a random swap argument to upper bound $\mathcal{D}^{2n}(R_{\epsilon,\gamma})$ with something that is independent of $\mathcal{D}$. We let $\Gamma_{n}$ denote the set of all permutations on $\{1,\dots,2n\}$ that potentially swap the $i$’th and $n+i$’th elements for all $i\in[n]$. So for all $\sigma\in\Gamma_{n}$ and $i\in[n]$, we have that either $\sigma(i)=i,\sigma(n+i)=n+i$ or $\sigma(i)=n+i,\sigma(n+i)=i$. From Lemma 3 in Attias23:regression (cf. also [Anthony_Bartlett_1999]), we get that for some $\pazocal{M}=\pazocal{Z}\times[K]$ and distribution $\mathcal{D}$ on $V$:
$\mathcal{D}^{2n}(R_{\epsilon,\gamma})=\mathop{\mathbb{E}}_{Z\sim\mathcal{D}^{2n}}\mathbb{P}_{\sigma\sim\text{Unif}(\Gamma_{n})}[\sigma(V)\in R_{\epsilon,\gamma}]\leq\max_{Z\in\pazocal{Z}^{2n}}\mathbb{P}_{\sigma\sim\text{Unif}(\Gamma_{n})}[\sigma(V)\in R_{\epsilon,\gamma}].$
The last step of the proof, requires reducing to learning a partial binary hypothesis class and using a disambiguation argument to turn the partial class into a total one, as in Attias23:regression. Concretely, we start by fixing a set $V\in\pazocal{M}^{2n}$, where $v_{i}=(\mathbf{z}_{i},h^{*}(\mathbf{z}_{i}))$ and a set $S=\{\mathbf{z}_{1},\dots,\mathbf{z}_{2n}\}$. For the projection, $\pazocal{H}(S)$ of the hypothesis class on the set, one can define the partial binary hypothesis class
@@
C.1 Learning without contexts
Suppose there is a fixed, unknown distribution, $\mathcal{D}$, over the set of possible follower types. We denote with $u(\mathbf{x},b_{f^{(i)}}(\mathbf{x}))$ the leader’s payoff when committing to a (possibly mixed) strategy $\mathbf{x}\in\Delta(\pazocal{A})$ against a follower of type $f^{(i)}$. We use the notation $u_{\mathcal{D}}(\mathbf{x})=\mathbb{E}_{f\sim\mathcal{D}}[u(\mathbf{x},b_{f}(\mathbf{x}))]$ and given a sample $S=\{f_{1},\dots,f_{m}\}$ of follower types, we write $u_{S}(\mathbf{x})=\frac{1}{m}\sum_{f\in S}u(\mathbf{x},b_{f}(\mathbf{x}))$.
Given such a sample, the goal of the leader is to choose a mixed strategy that fares well in expectation against any new sample from $\mathcal{D}$. Concretely, we want to choose $\hat{\mathbf{x}}\in\Delta(\pazocal{A})$ such that
$\mathbb{P}_{S}\left[\sup_{\mathbf{x}}u_{\mathcal{D}}(\mathbf{x})-u_{\mathcal{D}}(\hat{\mathbf{x}})>\epsilon\right]\leq\delta.$
Definition C.4 (Generalization guarantee).
A generalization guarantee for the leader in a Stackelberg game, $\pazocal{G}$, is a function $\epsilon_{\pazocal{G}}:\mathbb{Z}_{\geq 1}\times(0,1)\to\mathbb{R}_{\geq 0}$ such that for any $\delta\in(0,1)$, any $m\in\mathbb{Z}_{\geq 1}$, and any distribution $\mathcal{D}$ over the follower types, with probability at least $1-\delta$ over the draw of a set $S\sim\mathcal{D}^{m}$ for any $\mathbf{x}\in\Delta(\pazocal{A})$, the difference between the average payoff of $\mathbf{x}$ over $S$ and the expected payoff of $\mathbf{x}$ over $\mathcal{D}$ is at most $\epsilon_{\pazocal{G}}(m,\delta)$:
$\mathbb{P}_{S\sim\mathcal{D}^{m}}\left[\exists\mathbf{x}\in\Delta(\pazocal{A})\;\;\text{s.t.}\;\;|u_{\mathcal{D}}(\mathbf{x})-u_{S}(\mathbf{x})|>\epsilon_{\pazocal{G}}(m,\delta)\right]\leq\delta$
This implies a relationship between the utility of the leader’s strategy maximizing the average payoff over the training data to the expected utility of the optimal strategy. Formally, for $\hat{\mathbf{x}}\in\arg\sup_{\mathbf{x}\in\Delta(\pazocal{A})}u_{S}(\mathbf{x})$ and $\mathbf{x}^{*}\in\arg\sup_{\mathbf{x}\in\Delta(\pazocal{A})}u_{\mathcal{D}}(\mathbf{x})$
$\mathbb{P}_{S\sim\mathcal{D}^{m}}\left[|u_{\mathcal{D}}(\mathbf{x}^{*})-u_{S}(\hat{\mathbf{x}})|>2\epsilon_{\pazocal{G}}(m,\delta)\right]\leq\delta.$
We let $\pazocal{P}=\{u_{\mathbf{x}}:[K]\to\mathbb{R}\;\;\text{such that}\;\;u_{\mathbf{x}}(f)=u(\mathbf{x},b_{f}(\mathbf{x}))\}$ be the set of payoff functions for the leader. Each payoff function is parameterized by the leader’s mixed strategy, takes as input the follower’s type, and outputs the leader’s utility. We also define the dual class $\tilde{\pazocal{P}}=\{\tilde{u}_{f}:\Delta(\pazocal{A})\to\mathbb{R}\;\;\text{such that}\;\;\tilde{u}_{f}(\mathbf{x})=u(\mathbf{x},b_{f}(\mathbf{x}))\}$. Consider the following definition, introduced by balcan2018general and adapted to our setting.
Definition C.5.
A class $\pazocal{H}$ of hypotheses, $h:[K]\to\mathbb{R}$, is $(d,t)$-delineable if
  • For any $f\in[K]$, there is a set $\mathcal{J}$ of $t$ hyperplanes such that for any connected component $\pazocal{X}^{\prime}=\pazocal{X}\setminus\mathcal{J}$, the dual function $h_{f}(\mathbf{x})$ is linear over $\pazocal{X}^{\prime}$.
Lemma C.6.
Proof.
Next, we present the definition of pseudo-dimension [pollard2012convergence], which is complexity measure that characterizes a hypothesis class and in our setting helps us reason about how difficult the learning task of the leader is. Informally speaking, the more complex the hypothesis class, $\pazocal{P}$, the harder it is for the leader to distinguish between strategies, $\mathbf{x}$, which parameterize the class, and therefore, more samples are needed to learn.
Definition C.7 (Pseudo-dimension).
Using Lemma˜C.6 and results from the data-driven algorithm design literature [Balcan_2021, 10.1145/3676278, balcan2025generalization], we immediately get the following bound on the pseudo-dimension.
Theorem C.8 (Adaptation of Lemma 3.10 in balcan2018general).
It is known that the finiteness of the pseudo-dimension of a function class is a sufficient condition to show a generalization guarantee for the class [pollard2012convergence, dudley2010universal]. In our case, the above bound implies a generalization guarantee with function
Discussion. We note that the above generalization result is particularly strong. Informally, it asserts that for any distribution over the follower types, if the leader has so far encountered a number of followers that scales linearly with the size of her action set and logarithmically with the follower’s action set, it suffices for her to play a strategy that maximizes her average utility over the followers seen so far. This result holds even for an infinite number of follower types and extends prior work on learning in Stackelberg games without contextual information [letchford2009learning].
Remark C.9.
It is worth noting that even when there is only one context (i.e. $\pazocal{Z}$ is a singleton), our results in this section are incomparable with those in Section˜4. This is because in this section, the goal of the leader is to learn given a deterministic mapping from contexts to follower types. So, if the contexts are the same between all examples in the training set, the follower types corresponding to these examples will also be the same. In contrast, the context-free setting we consider in Section˜C.1 allows for a distribution over follower types. Thus the two bounds are generally incomparable.
Appendix D Computational complexity considerations
Computational complexity has been a longstanding concern in the analysis of Stackelberg games. In the simplest setting, where a leader interacts with a single known follower type conitzer2006computing have shown that computing the optimal mixed strategy for the leader can be achieved in polynomial time using linear programming. However, in more complex Stackelberg games such as Bayesian Stackelberg games, Stackelberg security games, and even 3-player Stackelberg games computing the leader’s optimal strategy has been shown to be NP-hard [conitzer2006computing, korzhyk10:stackelberg]. Therefore computational considerations have so far been largely ignored. In our setting, there are several interesting cases of structured Stackelberg games for which the leader’s optimal strategy can be learned in polynomial time with provable (albeit not optimal) regret guarantees.
In a structured Stackelberg game with hypothesis class $\pazocal{H}$ the leader’s optimal policy can be learned in polynomial time if $\pazocal{H}$ is efficiently online learnable. We give a high-level overview of this approach. First, we consider a learning algorithm for the hypothesis class $\pazocal{H}$. The leader can use this algorithm to predict the follower’s type and based on that compute the optimal mixed strategy against that type as if her prediction were correct. As shown by conitzer2006computing, for a fixed and known follower type, computing the leader’s optimal mixed strategy can be achieved in polynomial time using linear programming. The learner will make a finite number of follower type mispredictions before finding $h^{*}$, and at this point will play optimally against every follower type that appears in every subsequent round. Therefore, for any hypothesis class in the realizable setting that is efficiently online- (and therefore also PAC-) learnable we can establish a computationally efficient algorithm for learning the leader’s optimal policy. We now present one such example of a hypothesis class that can be learned efficiently when used to predict the followers’ types.
Example D.1.
Consider the hypothesis class, $\pazocal{H}_{\text{MDL}}$ of multiclass decision lists. Each decision list is a sequence of “if-then” rules, specifically of the form “if $\ell_{1}$ then $k_{1}$, else if $\ell_{2}$ then $k_{2}$, else if $\ell_{3}$ then $k_{3}$, …, else $k_{\text{final}}$”. Each decision list $h\in\pazocal{H}$ can be represented by a list $(\ell_{1},k_{1}),\ldots,(\ell_{n},k_{n}),(k_{\text{final}})]$ that consists of condition-value pairs $(\ell_{i},k_{i})$ and a final default value $(k_{\text{final}})$. Each condition $\ell_{i}$ is a literal (of either a variable in a set of input variables, $\pazocal{Z}=\{x_{1},\ldots,x_{n}\}$, or its negation) and each value $k_{i}$ belongs to a set of output labels, $[K]$. A decision list is evaluated sequentially. It takes as input an assignment of variables in $\pazocal{Z}$ i.e. $\mathbf{z}\in\{0,1\}^{n}$ and its output is the value corresponding to the first condition that is satisfied by the input, $\mathbf{z}$.
Theorem D.2.